OpenStudy (anonymous):

tan^(2)θ + 3secθ+3=0

OpenStudy (ash2326):

\[1+\tan^2 \theta=\sec^2 \theta\] So \[\tan^2\theta=\sec^2 \theta -1\] Put this in the given equation \[\sec^2\theta-1+3\sec \theta+3=0\] \[\sec^2\theta+3\sec \theta+2=0\] Put \[\sec \theta=x\] Now you have\[x^2+3x+2=0\] Solve for x, can you do that?

OpenStudy (anonymous):

Yep I just got it

OpenStudy (anonymous):

Finished it, just as you posted that. Tyvm tho :)

OpenStudy (anonymous):

My Final Answers were: cos = -1/2, cos = -1

OpenStudy (ash2326):

No problem :D

OpenStudy (abb0t):

What a clear explanation @ash2326

OpenStudy (anonymous):

Which converts to 120, 180, and 240 degrees. Got it =D

OpenStudy (anonymous):

Remember the identity: \[\sec^2\theta=\tan^2\theta+1\rightarrow \tan^2\theta=\sec^2\theta-1\] We re-write the equation as: \[(\sec^2(\theta) -1)+3\sec(\theta)+3=0\]\[\sec^2(\theta)+3\sec(\theta)+2=0\]This is a quadratic and so just factor it to get the possible values of sec(x). \[(\sec(\theta)+1)(\sec(\theta)+2)=0\] From the factorisation, either sec(x) + 1 = 0 or sec(x) + 2 = 0. Now we just solve for the two cases and get our values. Case 1: \[\sec(\theta)=-1\rightarrow \cos(\theta)=-1\]\[\theta=\pi\]Case 2:\[\sec(\theta)=-2\rightarrow \cos(\theta)=-\frac{1}{2}\]\[\theta=\frac{2\pi }{ 3 },\frac{ 4\pi }{ 3}\]\[\therefore \theta=\pi,\frac{ 2\pi }{ 3 },\frac{ 4\pi }{ 3 }\]